Net Present Value Method for a Service Company

Coast-to-Coast Inc. is considering the purchase of an additional delivery vehicle for $41,000 on January 1, 20Y1. The truck is expected to have a five-year life with an expected residual value of $5,000 at the end of five years. The expected additional revenues from the added delivery capacity are anticipated to be $75,000 per year for each of the next five years. A driver will cost $52,000 in 20Y1, with an expected annual salary increase of $4,000 for each year thereafter. The annual operating costs for the truck are estimated to be $3,000 per year.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

a. Determine the expected annual net cash flows from the delivery truck investment for 20Y1-20Y5.

Annual Net Cash Flow
20Y1 $
20Y2 $
20Y3 $
20Y4 $
20Y5 $

b. Compute the net present value of the investment, assuming that the minimum desired rate of return is 20%. Use the table of the present value of $1 presented above. When required, round to the nearest dollar. If required, use the minus sign to indicate a negative net present value.

Present value of annual net cash flows $
Investment
Net present value $

c. Is the additional truck a good investment based on your analysis?
, because the net present value is  
Answer #1

a. Determine the expected annual net cash flows from the delivery truck investment for 20Y1-20Y5.

Annual Net Cash Flow
20Y1 $75000-52000-3000-41000 = -21000
20Y2 $20000-4000 = 16000
20Y3 $16000-4000 = 12000
20Y4 $12000-4000 = 8000
20Y5 $8000-4000 = 4000+5000 = 9000

b. Compute the net present value of the investment, assuming that the minimum desired rate of return is 20%. Use the table of the present value of $1 presented above. When required, round to the nearest dollar. If required, use the minus sign to indicate a negative net present value.

Present value of annual net cash flows $42186
Investment -41000
Net present value $1186

c. Is the additional truck a good investment based on your analysis?
,Yes because the net present value is positive